# Conditional entropy and information of quantum processes

### v1

@@ IX Generalized mutual information of quantum channels
The mutual information $I(A;B)_{\rho}$ of a state $\rho_{AB}$ finds operational meaning in the task of erasing the total correlation present in the state $\rho_{AB}$ [80]. It also has direct operational meaning as the optimal error exponent in the asymmetric hypothesis testing, in Stein’s setting, for the state discrimination task between $\rho_{AB}$ (null hypothesis) and $\rho_{A}\otimes\rho_{B}$ (alternate hypothesis).
@@
For a multipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$, $\mathbb{I}[A;B;\ldots;N]_{\mathcal{N}}\geq 0$ and $\mathbb{I}^{\leftrightarrow}[A;B;\ldots;N]_{\mathcal{N}}\geq 0$ if the associated generalized divergence is always nonnegative for the states. When the associated generalized divergence is faithful for the states then $\mathbb{I}[A;B;\ldots;N]_{\mathcal{N}}=0$ and $\mathbb{I}^{\leftrightarrow}[A;B;\ldots;N]_{\mathcal{N}}=0$ if and only if $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$ is product of local channels, i.e., of the form $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}=\mathcal{N}^{1}_{A^{\prime}\to A}\otimes\mathcal{N}^{2}_{B^{\prime}\to B}\otimes\cdots\otimes\mathcal{N}^{n}_{N^{\prime}\to N}$.
@@
$I[A;B;\ldots;N]_{\mathcal{N}}=\inf_{\mathcal{M}^{1}\otimes\mathcal{M}^{2}\otimes\cdots\otimes\mathcal{M}^{n}}\mathbb{D}^{\leftrightarrow}[\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}\|\mathcal{M}^{1}_{A^{\prime}\to A}\otimes\mathcal{M}^{2}_{B^{\prime}\to B}\otimes\cdots\otimes\mathcal{M}^{n}_{N^{\prime}\to N}],$ (138)
@@
For a tele-covariant multipartite channel $\mathcal{T}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots B}$ [9, Theorem 5], the multipartite mutual information of the channel is equal to the multipartite mutual information of its Choi state $\Phi^{\mathcal{T}}_{R_{A}AR_{B}B\ldots R_{N}N}$,
@@
Proposition 10.
The multipartite mutual information of a multipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots B}$ is lower bounded by the multipartite mutual information of its Choi state $\Phi^{\mathcal{N}}_{R_{A}AR_{B}B\ldots R_{N}N}$,
$I[A;B;\ldots;N]_{\mathcal{N}}\geq I(R_{A}A;R_{B}B;\ldots;R_{N}N)_{\Phi^{\mathcal{N}}}.$ (144)
Proof.
$I[A;B;C]_{\mathcal{N}}$ $=D[\mathcal{N}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}\|\mathcal{M}^{1}_{A^{\prime}\to A}\otimes\mathcal{M}^{2}_{B^{\prime}\to B}\otimes\mathcal{M}^{3}_{C^{\prime}\to C}]$
$\geq D(\mathcal{N}(\Phi_{R_{A}A^{\prime}}\otimes\Phi_{R_{B}B^{\prime}}\otimes\Phi_{R_{C}C^{\prime}})\|\mathcal{M}^{1}(\Phi_{R_{A}A^{\prime}})\otimes\mathcal{M}^{2}(\Phi_{R_{B}B^{\prime}})\otimes\mathcal{M}^{3}(\Phi_{R_{C}C^{\prime}}))$
$=D(\Phi^{\mathcal{T}}_{R_{A}AR_{B}BR_{C}C}\|\Phi^{\mathcal{M}^{1}}_{R_{A}A}\otimes\Phi^{\mathcal{M}^{2}}_{R_{B}B}\otimes\Phi^{\mathcal{M}^{3}}_{R_{C}C})$
$\geq I(R_{A}A;R_{B}B;R_{C}C)_{\Phi^{\mathcal{T}}}.$ (145)
@@ IX.1 Conditional mutual information for channels: An alternate definition
Consider a tele-covariant tripartite channel $\mathcal{T}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}$. From Theorem 6, we have $I[A;C|B]_{\mathcal{T}}=I(R_{A}A;C|R_{B}BR_{C})_{\Phi^{\mathcal{T}}}$. We now notice that the reduced channel of a tele-covariant multipartite channel is also a tele-covariant channel and jointly tele-covariant with the full (original) channel, see Lemma 2 and [9, Remark 5]. Using this fact and Theorem 7,
@@
X Bidirectional conditional entropy
A bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is called a bidirectional channel if pairs $(A^{\prime},A)$ and $(B^{\prime},B)$ are in separate labs, say with Ash and Bob, respectively, and they have access only to the separable states for input to the channel (cf. [81, 72, 39]). However, it suffices to optimize over (pure) product states $\psi_{R_{A}A^{\prime}}\otimes\phi_{R_{B}B^{\prime}}$.
@@ X Bidirectional conditional entropy
$\mathbf{S}^{\leftrightarrow}[AB]_{\mathcal{N}}:=\mathbf{S}^{\leftrightarrow}[\mathcal{N}]:=-\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{R}_{B^{\prime}\to B}],$ (153)
@@ XI Discussion
The notions of entropy, conditional entropy, mutual information, and conditional mutual information form primitives of the information theory for both classical and quantum systems. The applications of the conditional entropy, the strong subadditivity of the entropy, and mutual information measures of the quantum states have been immense in the study of quantum systems and processes from fundamental as well as technological aspects, see, e.g., quantum error correction, quantum communication, quantum field theory, condensed matter physics, memory-effects in open quantum dynamics, quantum thermodynamics, etc. [82, 83, 84, 4, 8, 85, 86, 87, 88, 89, 74, 10].
@@
The fundamental role of conditional entropy and conditional mutual information in information theory motivates us to claim that these notions for quantum channels will have wide applications and implications in the study of quantum processes, including higher-order processes. We have discussed the recipe for the educated guesses of the definitions of the conditional entropy, strong subadditivity of the entropy, and the multipartite mutual information of quantum channels. This allows to extend the entropic and information measures of the channels to higher order processes in a way similar to [23, Section V and VI]. It would be useful to study the properties of these definitions and their operational meanings for transition maps (classical channels) [^3] To the best of our knowledge, we are not aware of these definitions for channels in classical information theory. We could perhaps be mistaken.. It is pertinent to utilize the conditional entropy of channels or their variants to algorithmically characterize their underlying causal structure [90, 91]. The conditional entropy of channel may also be used to study the entropic uncertainty relation for the measurements of quantum channels [92]. Our approach could be useful in the context of error correction in quantum gates (error correction at the dynamical level) and characterizing the memory effects in open quantum dynamics. We leave open the question of finding a novel information or computation task for which the conditional entropy of a quantum channel has exact operational meaning as the optimal rate.
@@ Acknowledgements.
The authors thank Sohail, Manabendra Nath Bera, Karol Horodecki, and Uttam Singh for the discussions. S.D. acknowledges support from the Science and Engineering Research Board, Department of Science and Technology (SERB-DST), Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also thanks IIIT Hyderabad for the Faculty Seed Grant. K.G. is supported by the Hong Kong Research Grant Council (RGC) through Grant No. 17307719 and 17307520.
Appendix A On generalized conditional entropies
Lemma 8.
The generalized conditional entropy is subadditive under tensor product, i.e., for a product of two bipartite channels $\mathcal{N}_{A^{\prime}C^{\prime}B^{\prime}D^{\prime}\to ACBD}:=\mathcal{N}^{1}_{A^{\prime}B^{\prime}\rightarrow AB}\otimes\mathcal{N}^{2}_{C^{\prime}D^{\prime}\rightarrow CD}$, we have
$\mathbf{S}[AC|BD]_{\mathcal{N}}\geq\mathbf{S}[A|B]_{\mathcal{N}^{1}}+\mathbf{S}[C|D]_{\mathcal{N}^{2}},$ (160)
if the associated generalized channel divergence $\mathbf{D}[\cdot\|\cdot]$ is additive under the tensor product of channels.
Proof.
The conditional entropy of $\mathcal{N}$ is given by
$-\mathbf{S}[AC|BD]_{\mathcal{N}}=\inf_{\mathcal{M}\in Q(B^{\prime}D^{\prime},BD)}\mathbf{D}\left[\mathcal{N}^{1}_{A^{\prime}B^{\prime}\rightarrow AB}\otimes\mathcal{N}^{2}_{C^{\prime}D^{\prime}\rightarrow CD}\|\mathcal{R}_{A^{\prime}C^{\prime}\rightarrow AC}\otimes\mathcal{M}_{B^{\prime}D^{\prime}\rightarrow BD}\right].$ (161)
The bipartite completely mixing map $\mathcal{R}_{A^{\prime}C^{\prime}\rightarrow AC}$ can be written as a product of two completely mixing maps, i.e., $\mathcal{R}_{A^{\prime}C^{\prime}\rightarrow AC}=\mathcal{R}_{A^{\prime}\rightarrow A}\otimes\mathcal{R}_{C^{\prime}\rightarrow C}$. The infimum is taken over the space of all bipartite channels $\mathcal{M}_{B^{\prime}D^{\prime}\rightarrow BD}$. If we only take the infimum over the space of product of channels $\mathcal{M}^{1}_{B^{\prime}\rightarrow B}\otimes\mathcal{M}^{2}_{D^{\prime}\rightarrow D}$ which is a subset of the whole space, then we have
$-\mathbf{S}[AC|BD]_{\mathcal{N}}$ $\leq\inf_{\mathcal{M}\in Q(B^{\prime},B)\otimes Q(D^{\prime},D)}\mathbf{D}\left[\mathcal{N}^{1}_{A^{\prime}B^{\prime}\rightarrow AB}\otimes\mathcal{N}^{2}_{C^{\prime}D^{\prime}\rightarrow CD}\|\mathcal{R}_{A^{\prime}\rightarrow A}\otimes\mathcal{R}_{C^{\prime}\rightarrow C}\otimes\mathcal{M}^{1}_{B^{\prime}\rightarrow B}\otimes\mathcal{M}^{2}_{D^{\prime}\rightarrow D}\right]$
$\leq\inf_{\mathcal{M}\in Q(B^{\prime},B)\otimes Q(D^{\prime},D)}\mathbf{D}\left[\mathcal{N}^{1}_{A^{\prime}B^{\prime}\rightarrow AB}\otimes\mathcal{N}^{2}_{C^{\prime}D^{\prime}\rightarrow CD}\|\mathcal{R}_{A^{\prime}\rightarrow A}\otimes\mathcal{M}^{1}_{B^{\prime}\rightarrow B}\otimes\mathcal{R}_{C^{\prime}\rightarrow C}\otimes\mathcal{M}^{2}_{D^{\prime}\rightarrow D}\right].$ (162)
If the generalized channel divergence $\mathbf{D}\left[\cdot\|\cdot\right]$ is additive under tensor product of channels, i.e, for any set of channels $\{\mathcal{N},\mathcal{M},\widetilde{\mathcal{N}},\widetilde{\mathcal{M}}\}$, it satisfies $\mathbf{D}\left[\mathcal{N}\otimes\widetilde{\mathcal{N}}\|\mathcal{M}\otimes\widetilde{\mathcal{M}}\right]=\mathbf{D}\left[\mathcal{N}\|\mathcal{M}\right]+\mathbf{D}\left[\widetilde{\mathcal{N}}\|\widetilde{\mathcal{M}}\right]$ (the geomtric Rényi divergence and max relative entropy are additive under tensor product of channels [34]), then we have $\mathbf{D}\left[\mathcal{N}^{1}_{A^{\prime}B^{\prime}\rightarrow AB}\otimes\mathcal{N}^{2}_{C^{\prime}D^{\prime}\rightarrow CD}\|\mathcal{R}_{A^{\prime}\rightarrow A}\otimes\mathcal{M}^{1}_{B^{\prime}\rightarrow B}\otimes\mathcal{R}_{C^{\prime}\rightarrow C}\otimes\mathcal{M}^{2}_{D^{\prime}\rightarrow D}\right]=\mathbf{D}\left[\mathcal{N}^{1}_{A^{\prime}B^{\prime}\rightarrow AB}\|\mathcal{R}_{A^{\prime}\rightarrow A}\otimes\mathcal{M}^{1}_{B^{\prime}\rightarrow B}\right]+\mathbf{D}\left[\mathcal{N}^{2}_{C^{\prime}D^{\prime}\rightarrow CD}\|\mathcal{R}_{C^{\prime}\rightarrow C}\otimes\mathcal{M}^{2}_{D^{\prime}\rightarrow D}\right]$. Using this in the above inequality, we obtain
$\mathbf{S}[AC|BD]_{\mathcal{N}}\geq\mathbf{S}[A|B]_{\mathcal{N}^{1}}+\mathbf{S}[C|D]_{\mathcal{N}^{2}}.$ (163)
Appendix B Proof of Theorem 1

### v2

@@ IX Generalized mutual information of quantum channels
The mutual information $I(A;B)_{\rho}$ of a state $\rho_{AB}$ finds operational meaning in the task of erasing the total correlation present in the state $\rho_{AB}$ [85]. It also has a direct operational meaning as the optimal error exponent in the asymmetric hypothesis testing, in Stein’s setting, for the state discrimination task between $\rho_{AB}$ (null hypothesis) and $\rho_{A}\otimes\rho_{B}$ (alternate hypothesis). Analogous to the mutual information of multipartite states, the multipartite generalized mutual information of a multipartite state $\rho_{AB\ldots N}$ is defined as
$\mathbb{I}(A;B;\ldots;N)_{\rho}:=\inf_{\omega_{i}\in{\mathrm{S}t}(i),\forall i\in\{A,B,\ldots,N\}}\mathbb{D}(\rho_{AB\ldots N}\|\otimes_{i=A}^{N}\omega_{i}).$ (152)
@@
For a multipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$, $\mathbb{I}[A;B;\ldots;N]_{\mathcal{N}}\geq 0$ if the associated generalized divergence is always nonnegative for the states. When the associated generalized divergence is faithful for the states then $\mathbb{I}[A;B;\ldots;N]_{\mathcal{N}}=0$ if and only if $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$ is product of local channels, i.e., of the form $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}=\mathcal{N}^{1}_{A^{\prime}\to A}\otimes\mathcal{N}^{2}_{B^{\prime}\to B}\otimes\cdots\otimes\mathcal{N}^{n}_{N^{\prime}\to N}$.
Proposition 10.
The multipartite generalized mutual information of a (multipartite) quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots B}$ is lower bounded by the multipartite generalized mutual information of its Choi state $\Phi^{\mathcal{N}}_{R_{A}AR_{B}B\ldots R_{N}N}$,
$\mathbb{I}[A;B;\ldots;N]_{\mathcal{N}}\geq\mathbb{I}(R_{A}A;R_{B}B;\ldots;R_{N}N)_{\Phi^{\mathcal{N}}}.$ (154)
Proof.
$\mathbb{I}[A;B;C]_{\mathcal{N}}$ $=\mathbb{D}[\mathcal{N}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}\|\mathcal{M}^{1}_{A^{\prime}\to A}\otimes\mathcal{M}^{2}_{B^{\prime}\to B}\otimes\mathcal{M}^{3}_{C^{\prime}\to C}]$
$\geq\mathbb{D}(\mathcal{N}(\Phi_{R_{A}A^{\prime}}\otimes\Phi_{R_{B}B^{\prime}}\otimes\Phi_{R_{C}C^{\prime}})\|\mathcal{M}^{1}(\Phi_{R_{A}A^{\prime}})\otimes\mathcal{M}^{2}(\Phi_{R_{B}B^{\prime}})\otimes\mathcal{M}^{3}(\Phi_{R_{C}C^{\prime}}))$
$=\mathbb{D}(\Phi^{\mathcal{N}}_{R_{A}AR_{B}BR_{C}C}\|\Phi^{\mathcal{M}^{1}}_{R_{A}A}\otimes\Phi^{\mathcal{M}^{2}}_{R_{B}B}\otimes\Phi^{\mathcal{M}^{3}}_{R_{C}C})$
$\geq\mathbb{I}(R_{A}A;R_{B}B;R_{C}C)_{\Phi^{\mathcal{N}}}.$ (155)
@@
$I[A;B;\ldots;N]_{\mathcal{N}}=\inf_{\mathcal{M}^{1}\otimes\mathcal{M}^{2}\otimes\cdots\otimes\mathcal{M}^{n}}D[\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}\|\mathcal{M}^{1}_{A^{\prime}\to A}\otimes\mathcal{M}^{2}_{B^{\prime}\to B}\otimes\cdots\otimes\mathcal{M}^{n}_{N^{\prime}\to N}],$ (156)
@@
For a tele-covariant multipartite channel $\mathcal{T}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots B}$ [10, Theorem 5], the multipartite mutual information of the channel is equal to the multipartite mutual information of its Choi state $\Phi^{\mathcal{T}}_{R_{A}AR_{B}B\ldots R_{N}N}$,
@@
A direct consequence of the above theorem is the following corollary.
Corollary 7.
The half of the mutual information of any tele-covariant bipartite unitary channel $\mathcal{U}_{A^{\prime}B^{\prime}\to AB}$ is equal to the von Neumann entropy of its reduced state,
$\frac{1}{2}I[A;B]_{\mathcal{U}}=S(R_{A}A)_{\Phi^{\mathcal{U}}}=S(R_{B}B)_{\Phi^{\mathcal{U}}}.$ (162)
Lemma 6.
The multipartite max-mutual information and the multipartite geometric Rényi mutual information of a (multipartite) quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$ are respectively given as
$I_{\max}[A;B;\ldots;N]_{\mathcal{N}}$ $=\inf_{\bigotimes_{i=A}^{N}\mathcal{M}^{i}}D_{\max}(\Phi^{\mathcal{N}}_{R_{A}AR_{B}B\ldots R_{N}N}\|\otimes_{i=A}^{N}\Phi^{\mathcal{M}^{i}}_{R_{i}i}),$ (163)
$\widehat{I}_{\alpha}[A;B;\ldots;N]_{\mathcal{N}}$ $=\inf_{\bigotimes_{i=A}^{N}\mathcal{M}^{i}}\widehat{D}_{\alpha}(\Phi^{\mathcal{N}}_{R_{A}AR_{B}B\ldots R_{N}N}\|\otimes_{i=A}^{N}\Phi^{\mathcal{M}^{i}}_{R_{i}i})\quad\forall\alpha\in(1,2],$ (164)
where $\mathcal{M}^{i}\in{\mathrm{Q}}(i^{\prime},i)$ for all $i\in\{A,B,\ldots,N\}$.
@@ IX.1 Conditional mutual information for channels: An alternate definition
Consider a tele-covariant tripartite channel $\mathcal{T}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}$. From Theorem 6, we have $I[A;C|B]_{\mathcal{T}}=I(R_{A}A;C|R_{B}BR_{C})_{\Phi^{\mathcal{T}}}$. We now notice that the reduced channel of a tele-covariant multipartite channel is also a tele-covariant channel and jointly tele-covariant with the full (original) channel, see Lemma 2 and [10, Remark 5]. Using this fact and Theorem 7,
@@
X Bidirectional conditional entropy and information
A bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is called a bidirectional channel if pairs $(A^{\prime},A)$ and $(B^{\prime},B)$ are in separate labs, say with Ash and Bob, respectively, and they have access only to the separable states for input to the channel (cf. [86, 77, 42]). However, it suffices to optimize $\mathbf{D}^{\leftrightarrow}[\mathcal{N}^{1}\|\mathcal{N}^{2}]$ between two bidirectional quantum channels $\mathcal{N}^{1},\mathcal{N}^{2}\in{\mathrm{Q}}(A^{\prime}B^{\prime},AB)$ over (pure) product input states $\psi_{R_{A}A^{\prime}}\otimes\phi_{R_{B}B^{\prime}}$.
@@ X Bidirectional conditional entropy and information
$\mathbf{S}^{\leftrightarrow}[AB]_{\mathcal{N}}:=\mathbf{S}^{\leftrightarrow}[\mathcal{N}]:=-\mathbf{D}^{\leftrightarrow}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{R}_{B^{\prime}\to B}],$ (172)
@@
Definition 19.
Lemma 9.
For a multipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$, $\mathbb{I}^{\leftrightarrow}[A;B;\ldots;N]_{\mathcal{N}}\geq 0$ if the associated generalized divergence is always nonnegative for the states. When the associated generalized divergence is faithful for the states then $\mathbb{I}^{\leftrightarrow}[A;B;\ldots;N]_{\mathcal{N}}=0$ if and only if $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}$ is product of local channels, i.e., of the form $\mathcal{N}_{A^{\prime}B^{\prime}\ldots N^{\prime}\to AB\ldots N}=\mathcal{N}^{1}_{A^{\prime}\to A}\otimes\mathcal{N}^{2}_{B^{\prime}\to B}\otimes\cdots\otimes\mathcal{N}^{n}_{N^{\prime}\to N}$.
@@ XI Discussion
The notions of entropy, conditional entropy, mutual information, and conditional mutual information form primitives of the information theory for both classical and quantum systems. The applications of the conditional entropy, the strong subadditivity of the entropy, and mutual information measures of the quantum states have been immense in the study of quantum systems and processes from fundamental as well as technological aspects, see, e.g., quantum error correction, quantum communication, quantum field theory, condensed matter physics, memory-effects in open quantum dynamics, quantum thermodynamics, etc. [87, 88, 89, 5, 90, 9, 91, 92, 93, 94, 31, 95, 79, 11].
@@
The fundamental role of conditional entropy and conditional mutual information in information theory motivates us to claim that these notions for quantum channels will have wide applications and implications in the study of quantum processes, including higher-order processes. We have discussed the recipe for the educated guesses of the definitions of the conditional entropy, strong subadditivity of the entropy, and the multipartite mutual information of quantum channels. This allows to extend the entropic and information measures of the channels to higher order processes in a way similar to [24, Section V and VI]. It would be useful to study the properties of these definitions and their operational meanings for transition maps (classical channels) [^3] To the best of our knowledge, we are not aware of definitions of the conditional entropies and (conditional) mutual information for multipartite channels in classical information theory. We could perhaps be mistaken.. One may also utilize the conditional entropy of channels or their variants to algorithmically characterize their underlying causal structure [96, 97] and characterize the memory effects in open quantum dynamics [98, 11]. The conditional entropy of channel may also be used to derive the entropic uncertainty relation for the measurements of quantum channels [99]. Our approach could be useful in the context of error correction in quantum gates (error correction at the dynamical level) and characterizing the memory effects in open quantum dynamics. We leave open the question of finding a novel information or computation task, other than the channel discrimination, for which the conditional entropy of an arbitrary quantum channel has operational meaning as the optimal rate.
@@ Acknowledgements.
The authors thank Sohail, Himanshu Badhani, Manabendra Nath Bera, Karol Horodecki, and Uttam Singh for the discussions. S.D. acknowledges support from the Science and Engineering Research Board, Department of Science and Technology (SERB-DST), Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also thanks IIIT Hyderabad for the Faculty Seed Grant. K.G. is supported by the Senior Research Fellowship Scheme SRFS2021-7S02 and the John Templeton Foundation through grant 62312, “The Quantum Information Structure of Spacetime” (qiss.fr).
Appendix A Proof of Theorem 1